If one dimension of a square is increased by 3cm and the other dimension is decreased by 2cm, the perimeter of the resulting rectangle is 120% of the perimeter of the square. Find the length of one side of the square
Answer by richwmiller(17219)

2*(x+3)+2*(x+2)=1.2*(4x)
divide by 2
x+3+x+2=1.2*2x
2x+5=2.4x
5=.4x
x=12.5